OpenStudy (anonymous):

Rosa works due north of home. Her husband Juan works due east. They leave for work at the same time. By the time Rosa is 5 miles from home, the distance between them is one more mile than Juan's distance from home. How far from home is Juan?

OpenStudy (radar):

 Pythagaroem \[25+x ^{2}=(x+1)^{2}\] \[25+x ^{2}=x ^{2}+2x+1\] 2x=24 x=12 Mr. Juan is 12 miles out. (excuse spelling of Pyth...........